Solve the given Bernoulli equation by using this substitution. t2y' + 2ty − y3 = 0, t > 0
Answer #1


Given Bernoullis equation : yond y + @ y 2 - Let us take 1 = y-e Thus from ② +30= v=- —B) = e- 4 int e inte = 141. Thus multiplying (3) by it ( ) - Thus, la ley ) = -24-6 dt - Where te is an integration constant +-6+1 + C - -6+1 + +4 = t